6. A subway train is travelling at 80km/h when it approaches a slower train 50 m ahead
travelling in the same direction at 25 km/h. if the faster train begins to decelerate at 2.6
m/s²
. While the slower train continues at constant speed. How soon and what relative
speed will they collide.
1
Expert's answer
2020-05-03T17:49:04-0400

"V_1=80\;km/h=80\frac{1000}{3600}=22.22 \;m/s."

"V_2=25\;km/h=25\frac{1000}{3600}=6.94 \;m/s."

"V_2t+50=V_1t-\frac{at^2}{2}" .

"6.94t+50=22.22t-\frac{2.6t^2}{2}."

"1.3t^2-15.28t+50=0."

This equation does not have real roots.

Therefore, the trains never collide.